OpenStudy (lexluth0r):

Can someone help guide me through this? Solve the system: y=x(squared)-3x-8 y=x+4

OpenStudy (phi):

you can show x squared like this: x^2

OpenStudy (lexluth0r):

Oh ok thanks I'll remember that

OpenStudy (dayakar):

y = x^2-3x-8--------(1) y= x+4 ---------(2) (2) = (1) [ two equations right hand sides are equal] x^2-3x-8 = x+4 [ L. H.S ARE EQUAL] SOLVE FOR x

OpenStudy (lexluth0r):

Ok I combined them and got x^2-4x-12=0 What method should I use to solve for x

OpenStudy (phi):

You can factor that (if you know how)